# LLM2Vec: Large Language Models Are Secretly Powerful Text Encoders

### v1

@@ Abstract
Large decoder-only language models (LLMs) are the state-of-the-art models on most of today’s NLP tasks and benchmarks. Yet, the community is only slowly adopting these models for text embedding tasks, which require rich contextualized representations. In this work, we introduce LLM2Vec, a simple unsupervised approach that can transform any decoder-only LLM into a strong text encoder. LLM2Vec consists of three simple steps: 1) enabling bidirectional attention, 2) masked next token prediction, and 3) unsupervised contrastive learning. We demonstrate the effectiveness of LLM2Vec by applying it to 3 popular LLMs ranging from 1.3B to 7B parameters and evaluate the transformed models on English word- and sequence-level tasks. We outperform encoder-only models by a large margin on word-level tasks and reach a new unsupervised state-of-the-art performance on the Massive Text Embeddings Benchmark (MTEB). Moreover, when combining LLM2Vec with supervised contrastive learning, we achieve state-of-the-art performance on MTEB among models that train only on publicly available data. Our strong empirical results and extensive analysis demonstrate that LLMs can be effectively transformed into universal text encoders in a parameter-efficient manner without the need for expensive adaptation or synthetic GPT-4 generated data.
@@ 1 Introduction
Text embedding models aim to encode the semantic content of natural language text in vector representations which then facilitate various natural language processing (NLP) tasks, such as semantic textual similarity, information retrieval, and clustering. For many years, the dominating paradigm for building such models relied on pre-trained bidirectional encoders or encoder-decoders such as BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020), which are typically adapted for text embedding tasks by following a multi-step training pipeline consisting of weakly- and fully-supervised contrastive training (Ni et al., 2022; Li et al., 2023; Xiao et al., 2023, inter alia). Only recently, the community started to adopt decoder-only LLMs for embedding text (Muennighoff, 2022; Ma et al., 2023; Wang et al., 2023; Springer et al., 2024).
@@
We apply LLM2vec to 3 decoder-only LLMs ranging from 1.3B to 7B parameters (S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B) and evaluate the resulting models on word- and sequence-level tasks. On word-level tasks (chunking, named-entity recognition, and part-of-speech tagging), LLM2Vec-transformed models outperform strong encoder-only models by a large margin, demonstrating its effectiveness for producing rich contextualized token representations. On the Massive Text Embeddings Benchmark (MTEB), LLM2Vec-transformed models set a new state-of-the-art for unsupervised models, with our best model reaching a score of $56.8$. Additionally, we combine LLM2Vec with supervised contrastive training and achieve a new state-of-the-art performance among models that train only on publicly available data. Beyond our strong empirical results, we provide an extensive analysis of how LLM2Vec affects the representations of the underlying model and reveal an intriguing property of Mistral-7B, the currently best performing architecture on MTEB.
Overall, our work demonstrates that decoder-only LLMs are indeed capable of producing universal text embeddings and only very little adaptation is required to reveal this ability. Our code and pre-trained models are available at https://github.com/McGill-NLP/llm2vec.
@@ 2.1 Three simple ingredients
Enabling bidirectional attention (Bi)
The first step of the LLM2Vec approach is to replace the causal attention mask of decoder-only LLMs by an all-ones matrix (see Section B.1 for background on the self-attention mechanism). This gives each token access to every other token in the sequence, converting it into a bidirectional LLM (Devlin et al., 2019; Liu et al., 2019). However, it is not a priori clear why this should lead to better sequence representations. After all, the decoder-only LLM was not trained to attend to future tokens and therefore, this naive approach might even lead to worse representations. As we show, simply enabling bidirectional attention does indeed decrease embedding performance for most models. We can however easily adapt a model to make use of its bidirectional attention.
Masked next token prediction (MNTP)
We propose a simple strategy to make the model aware of its bidirectional attention by adapting it via masked next token prediction (MNTP). MNTP is a novel training objective that combines next token prediction with masked language modeling (Devlin et al., 2019). Given an arbitrary sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{N})$ as input, we first mask a fraction of the input tokens and then train the model to predict the masked tokens based on the past and future context. Crucially, when predicting a masked token at position $i$, we compute the loss based on the logits obtained from the token representation at the previous position $i-1$, not the masked position itself (see Figure 1).
Unsupervised contrastive learning (SimCSE)
While the previous two steps of the LLM2Vec recipe can transform any decoder-only LLM into an encoder for word-level tasks, they might not be sufficient for sequence representations. Unlike bidirectional encoders that include a next sentence prediction objective in their pre-training objectives (Devlin et al., 2019), decoder-only LLMs are not explicitly trained to capture the context of the entire sequence. To fill this gap, we apply unsupervised contrastive learning via SimCSE (Gao et al., 2021). Specifically, given an input sequence, it is passed through the model twice with independently sampled dropout masks, resulting in two different representations for the same sequence. The model is trained to maximize the similarity between these two representations while minimizing the similarity with representations of other sequence in the batch. Crucially, this step does not require any paired data and can be applied using any collection of sequences. We use a pooling operation on the token representations to get the sequence representation (more details in Section 3.2).
The three steps combined – Bi + MNTP + SimCSE – constitute the LLM2Vec recipe.
@@ 2.2 Transforming decoder-only LLMs with LLM2Vec
We experiment with 3 different decoder-only LLMs ranging from 1.3B to 7B parameters: Sheared-LLaMA-1.3B (S-LLaMA-1.3B, Xia et al., 2023), Llama-2-7B-chat (LLaMA-2-7B, Touvron et al., 2023), and Mistral-7B-Instruct-v0.2 (Mistral-7B, Jiang et al., 2023).
@@
We perform both the MNTP and the unsupervised SimCSE step using data from English Wikipedia. We select data from Wikipedia as it was included in the pre-training mixture of all the models we experiment with. It is therefore fair to assume that these two adaptation steps are not teaching the model any new knowledge beyond how to attend to future tokens and how to construct sequence representations. Specifically, we use the Wikitext-103 dataset (Merity et al., 2017) for the MNTP step and a subset of Wikipedia sentences released by Gao et al. (2021) for the unsupervised SimCSE step.
@@
We follow established practice from the masked language modeling literature and randomly mask a fraction of the tokens from the input sequence (Devlin et al., 2019; Liu et al., 2019). We use the underscore (_) as the mask token, since the models we experiment with do not have a special token for masking. We fine-tune the model using LoRA (Hu et al., 2022) to predict the masked token using the representation of the previous token to maximally align our training objective with the pre-training setup of decoder-only LLMs. For all models, we trained for 1000 steps with a batch size of 32 on a single 80GB A100 GPU. For 7B models, this training takes only 90 minutes. We provide additional details of our training setup and hyperparameters in Section D.1.1.
@@
For the contrastive training, we apply the unsupervised SimCSE approach by Gao et al. (2021). The positive examples are constructed by applying the LLM’s dropout twice on the same input sequence, whereas other sequences in the batch act as in-batch negatives. We merge the MNTP LoRA weights into the base model and initialize new LoRA parameters before starting the SimCSE training, which ensures that the model retains the knowledge learned in the previous step. Similar to the MNTP step, we train for 1000 steps. For 7B models, this training takes 2.5 hours on a single 80GB A100 GPU with a batch size of 128. We provide additional details of our training setup and hyperparameters in Section D.1.2.
@@ 3.1 Evaluation on word-level tasks
We evaluate on three word-level tasks: chunking, named-entity recognition (NER), and part-of-speech tagging (POS), using the CoNLL-2003 benchmark (Tjong Kim Sang & De Meulder, 2003). We embed each input sentence and train a task-specific linear classifier on top of the frozen representations. This is akin to the linear probing setup commonly used in the language model analysis literature (Belinkov, 2022). We compare the LLM2Vec-transformed models to DeBERTa-v3-large (He et al., 2023), the current state-of-the-art encoder-only model. Additional details about our setup are provided in Section D.1.3.
@@ Results
Next, we observe that for all models and tasks, further adapting via MNTP (i.e. LLM2Vec w/o SimCSE) improves performance. For instance, in the chunking task, we see improvements for S-LLaMA-1.3B (by $5\%$), LLaMA-2-7B (by $4\%$), and Mistral-7B (by $4\%$). Finally, combining MNTP with SimCSE, however, performs worse than just applying MNTP. This is expected for word-level tasks, as SimCSE adapts the representations for sequence-level tasks.
@@ 3.2 Evaluation on sequence-level tasks
Figure 3: Unsupervised results on our 15-task subset of the MTEB dataset. We ablate three different pooling choices: EOS, mean pooling, and weighted mean pooling. LLM2Vec is compatible with all three approaches and works best with mean pooling.
@@
Echo embeddings 22.68 51.07 36.78 75.87 72.69 73.60 29.54 50.26
Table 1: Unsupervised results on MTEB. We compare S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B with and without LLM2Vec to the unsupervised BERT models of Gao et al. (2021) as well as Echo embeddings (Springer et al., 2024).
Next, we evaluate on the Massive Text Embedding Benchmark (MTEB), a collection of 7 diverse embedding task categories covering a total of 56 datasets (Muennighoff et al., 2023). To select the best-performing pooling method for each method, we perform ablations on a 15-task subset consisting of representative tasks from each of the MTEB categories. We provide additional details and justification for how we chose this subset in Section C.1.
Setup
Following previous work (Su et al., 2023; Wang et al., 2023; Springer et al., 2024), we evaluate with task-specific instructions. We use the same instructions as Wang et al. (2023) and only add instructions to queries. The instructions can be found in Table 10. For symmetric tasks, the same instruction will be used for the query and the document. When applying (weighted) mean pooling (Muennighoff, 2022), we exclude the instruction tokens.
As a baseline, we compare to the unsupervised BERT models obtained from Gao et al. (2021). Additionally, we compare to Echo embeddings, a concurrent approach by Springer et al. (2024), which we run with the same models and instructions (see Section E.1 for more details on our implementation of Echo embeddings). Echo duplicates the input and pools only over the second occurrence to address the limitation of causal attention.
Results on our 15-task subset of MTEB
Figure 3 shows the impact of various pooling methods for all three models on the subset of MTEB tasks. We can clearly observe that applying causal attention is sub-optimal when constructing text embeddings. The naive approach of applying the EOS pooling for models with causal attention is outperformed by (weighted) mean pooling. Enabling bidirectional attention without any training harms performance for S-LLaMA-1.3B and LLaMA-2-7B. Similar to our word-level results, the performance of Mistral-7B improves with bidirectional attention, even without any training.
For LLM2Vec-transformed models, applying MNTP training improves the performance of S-LLaMA-1.3B and Mistral-7B. Moreover, applying SimCSE further boosts the performance of S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B by $49.8\%$, $23.2\%$, and $37.5\%$ compared to the best causal baseline on the MTEB subset. We conduct further ablations of each component of the LLM2Vec recipe in Section D.2.2 (Table 5).
@@
Table 1 shows the results of the best-performing models, which we select based on the ablation above, on the full MTEB dataset. After the first two steps of LLM2Vec—bidirectional attention and MNTP (LLM2Vec w/o SimCSE)—we observe a considerable improvement in performance for all three models (e.g., $16.4\%$ improvement for Mistral-7B).
When comparing to Echo embeddings, LLM2Vec (w/o SimCSE) [^3] We only directly compare the performance after the first two steps of LLM2Vec to Echo embeddings as applying SimCSE involves learning sequence representation, which makes the comparison unfair. leads to improved performance for S-LLaMA-1.3B and LLaMA-2-7B and performs almost on par for Mistral-7B. However, compared to Echo embeddings, LLM2Vec is much more efficient as Echo embeddings repeat the input and therefore double the sequence length which makes inference considerably slower (we provide a runtime comparison in Section E.2).
Adding the final step of the LLM2Vec recipe— unsupervised SimCSE—further boosts all three models by a large margin, making our LLM2Vec Mistral-7B state-of-the-art among all unsupervised models with a score of $56.80$.
Overall, our results highlight that LLM2Vec is successful at transforming decoder-only LLMs into strong text embedding models that outperform previous unsupervised approaches on the challenging MTEB leaderboard.

### v2

@@ Abstract
Large decoder-only language models (LLMs) are the state-of-the-art models on most of today’s NLP tasks and benchmarks. Yet, the community is only slowly adopting these models for text embedding tasks, which require rich contextualized representations. In this work, we introduce LLM2Vec, a simple unsupervised approach that can transform any decoder-only LLM into a strong text encoder. LLM2Vec consists of three simple steps: 1) enabling bidirectional attention, 2) masked next token prediction, and 3) unsupervised contrastive learning. We demonstrate the effectiveness of LLM2Vec by applying it to 4 popular LLMs ranging from 1.3B to 8B parameters and evaluate the transformed models on English word- and sequence-level tasks. We outperform encoder-only models by a large margin on word-level tasks and reach a new unsupervised state-of-the-art performance on the Massive Text Embeddings Benchmark (MTEB). Moreover, when combining LLM2Vec with supervised contrastive learning, we achieve state-of-the-art performance on MTEB among models that train only on publicly available data (as of May 24, 2024). Our strong empirical results and extensive analysis demonstrate that LLMs can be effectively transformed into universal text encoders in a parameter-efficient manner without the need for expensive adaptation or synthetic GPT-4 generated data.
@@ 1 Introduction
Text embedding models aim to encode the semantic content of natural language text in vector representations which then facilitate various natural language processing (NLP) tasks, such as semantic textual similarity, information retrieval, and clustering. For many years, the dominating paradigm for building such models relied on pre-trained bidirectional encoders or encoder-decoders such as BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020), which are typically adapted for text embedding tasks by following a multi-step training pipeline consisting of weakly- and fully-supervised contrastive training (Ni et al., 2022; Li et al., 2023a; Xiao et al., 2023, inter alia). Only recently, the community started to adopt decoder-only LLMs for embedding text (Muennighoff, 2022; Ma et al., 2023; Wang et al., 2023; Springer et al., 2024; Li & Li, 2024).
@@
We apply LLM2vec to 4 decoder-only LLMs ranging from 1.3B to 8B parameters (S-LLaMA-1.3B, LLaMA-2-7B, Mistral-7B, Meta-LLaMA-3-8B) and evaluate the resulting models on word- and sequence-level tasks. On word-level tasks (chunking, named-entity recognition, and part-of-speech tagging), LLM2Vec-transformed models outperform strong encoder-only models by a large margin, demonstrating its effectiveness for producing rich contextualized token representations. On the Massive Text Embeddings Benchmark (MTEB), LLM2Vec-transformed models set a new state-of-the-art for unsupervised models, with our best model reaching a score of $56.8$. Additionally, we combine LLM2Vec with supervised contrastive training and achieve a new state-of-the-art performance among models that train only on publicly available data. Beyond our strong empirical results, we provide an extensive analysis of how LLM2Vec affects the representations of the underlying model and reveal an intriguing property of Mistral-7B, which can handle bidirectional attention without any fine-tuning.
Overall, our work demonstrates that decoder-only LLMs are indeed capable of producing universal text embedding and only very little adaptation is required to reveal this ability. Our code and pre-trained models is publicly available at https://github.com/McGill-NLP/llm2vec.
@@ 2.1 Three simple ingredients
Enabling bidirectional attention
The first step of the LLM2Vec approach is to replace the causal attention mask of decoder-only LLMs by an all-ones matrix (see Section B.1 for background on the self-attention). This gives each token access to every other token in the sequence, converting it into a bidirectional LLM (Devlin et al., 2019; Liu et al., 2019). However, it is not a priori clear, why this should lead to better sequence representations. After all, the decoder-only LLM was not trained to attend to future tokens and therefore, this naive approach might even lead to worse representations. As we show, simply enabling bidirectional attention does indeed decrease in embedding performance for most models. We can however easily adapt a model to make use of its bidirectional attention.
Masked next token prediction
We use a simple strategy to make the model aware of its bidirectional attention by adapting it via masked next token prediction (MNTP). MNTP is a training objective that combines next token prediction with masked language modeling (Lv et al., 2023). Given an arbitrary sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{N})$ as input, we first mask a fraction of the input tokens and then train the model to predict the masked tokens based on the past and future context. Crucially, when predicting a masked token at position $i$, we compute the loss based on the logits obtained from the token representation at the previous position $i-1$, not the masked position itself (see Figure 1).
Unsupervised contrastive learning
While the previous two steps of the LLM2Vec recipe can transform any decoder-only LLM into an encoder for word-level tasks, they might not be sufficient for sequence representations. Unlike bidirectional encoders that include a next sentence prediction objective in their pre-training objectives (Devlin et al., 2019), decoder-only LLMs are not explicitly trained to capture the context of the entire sequence. To fill this gap, we apply unsupervised contrastive learning via SimCSE (Gao et al., 2021). Specifically, given an input sentence, it is passed through the model twice with independently sampled dropout masks, resulting in two different representations for the same sentence. The model is trained to maximize the similarity between these two representations while minimizing the similarity with representations of other sentences in the batch. Crucially, this step does not require any sentence pair data and can be applied using any collection of sentences. We use a pooling operation on the word representations to get the sentence representation (more details in Section 3.2).
@@ 2.2 Transforming decoder-only LLMs with LLM2Vec
For most of our results, we experiment with 3 different decoder-only LLMs ranging from 1.3B to 7B parameters: Sheared-LLaMA-1.3B (S-LLaMA-1.3B, Xia et al., 2023), Llama-2-7B-chat (LLaMA-2-7B, Touvron et al., 2023), and Mistral-7B-Instruct-v0.2 (Mistral-7B, Jiang et al., 2023a). In Tables 1 and 2, we provide additional results for the recently released Meta-Llama-3-8B-Instruct model (Meta-LLaMA-3-8B, AI@Meta, 2024).
@@
We perform both the MNTP and the unsupervised SimCSE step using data from English Wikipedia. We select data from Wikipedia as it is presumably included in the pre-training mixture of all the models we experiment with. It is therefore fair to assume that these two adaptation steps are not teaching the model any new knowledge beyond how to attend to future tokens and how to construct sequence representations. Specifically, we use the Wikitext-103 dataset (Merity et al., 2017) for the MNTP step and a subset of Wikipedia sentences released by Gao et al. (2021) for the unsupervised SimCSE step.
@@
We follow established practice from the masked language modeling literature and randomly mask a fraction of the tokens from the input sequence (Devlin et al., 2019; Liu et al., 2019). We use the underscore (_) as the mask token, since the models we experiment with do not have a special token for masking. We fine-tune the model using LoRA (Hu et al., 2022) to predict the masked token using the representation of the previous token to maximally align our training objective with the pre-training setup of decoder-only LLMs. For all models, we trained for 1000 steps with a batch size of 32 on a single 80GB A100 GPU. For 7B and 8B models, this training takes only 100 minutes. We provide additional details of our training setup and hyperparameters in Section D.1.1.
@@
For the contrastive training, we apply the unsupervised SimCSE approach by Gao et al. (2021). The positive examples are constructed by applying LLM’s dropout twice on the same input sequence, whereas the other sequences in the batch act as in-batch negatives. We merge the MNTP LoRA weights into the base model and initialize new LoRA parameters before starting the SimCSE training, which ensures that the models retains the knowledge learned in the previous step. Similar to the MNTP step, we train for 1000 steps. For 7B and 8B models, this training takes 3 hours on a single 80GB A100 GPU with a batch size of 128. We provide additional details of our training setup and hyperparameters in Section D.1.2.
@@ 3.1 Evaluation on word-level tasks
We evaluate three word-level tasks: chunking, named-entity recognition (NER), and part-of-speech tagging (POS), using the CoNLL-2003 benchmark (Tjong Kim Sang & De Meulder, 2003). We embed each input sentence and train a task-specific linear classifier on top of the frozen representations. This is akin to the linear probing setup commonly used in the language model analysis literature (Belinkov, 2022). We compare the LLM2Vec-transformed models to DeBERTa-v3-large (He et al., 2023), the current state-of-the-art encoder-only model. Additional details about our setup are provided in Section D.1.3.
@@ Results
Focusing on the LLM2Vec-transformed models, we observe that for all models and tasks, adapting via MNTP improves performance. For instance, in the chunking task, we see improvements for S-LLaMA-1.3B (by $5\%$), LLaMA-2-7B (by $4\%$), and Mistral-7B (by $4\%$). Combining MNTP with SimCSE, however, performs worse than just applying MNTP. This is expected for word-level tasks, as SimCSE adapts the representations for sequence-level tasks.
@@ 3.2 Evaluation on sequence-level tasks
Figure 3: Unsupervised results on our 15 task subset of the MTEB dataset. We ablate three different pooling choices: EOS, mean pooling, and weighted mean pooling. LLM2Vec is compatible with all three approaches and works best with mean pooling.
Next, we evaluate on the Massive Text Embedding Benchmark (MTEB), a collection of 7 diverse embedding task categories covering a total of 56 datasets (Muennighoff et al., 2023). To select the best-performing pooling method for each method, we perform ablations on a 15 task subset consisting of representative tasks from each of the MTEB categories. We provide additional details and justification for how we chose this subset in Section C.1.
Setup
Following previous work (Su et al., 2023; Wang et al., 2023; Springer et al., 2024), we evaluate with task-specific instructions. For a fair comparison, we use the same set of instructions as Wang et al. (2023) which are also used by Springer et al. (2024). The instructions are only added to queries and can be found in Table 10 of Section C.2. For symmetric tasks, the same instruction will be used for the query and the document. When applying (weighted) mean pooling (Muennighoff, 2022), we exclude the instruction tokens.
As a baseline, we compare to the unsupervised BERT models obtained from Gao et al. (2021). Additionally, we compare to Echo embeddings, a concurrent approach by Springer et al. (2024), which we run with the same models and instructions (see Section E.1 for more details on our implementation of Echo embeddings). Echo duplicates the input and takes the pooling over the second occurrence to address the limitation of causal information flow.
Results on our 15 task subset of MTEB
Figure 3 shows the impact of various pooling methods for all three models on the subset of MTEB tasks. We can clearly observe that applying causal attention is sub-optimal when constructing text embeddings. The dominant paradigm of applying the EOS pooling for models with causal attention is outperformed by (weighted) mean pooling. Enabling bidirectional attention without any training harms performance for S-LLaMA-1.3B and LLaMA-2-7B. Similar to our word-level results, the performance of Mistral-7B improves with bidirectional attention, even without any training.
For LLM2Vec-transformed models, applying MNTP training improves the performance of S-LLaMA-1.3B and Mistral-7B. Moreover, applying SimCSE further boosts the performance of S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B by $49.8\%$, $23.2\%$, and $37.5\%$ compared to the best causal baseline on the MTEB subset. We further conduct an ablation of each component of LLM2Vec recipe in Section D.2.2 (Table 5).
@@
Echo 22.68 51.07 36.78 75.87 72.69 73.60 29.54 50.26
Meta-LLaMA-3-8B
Uni + w. Mean 15.17 46.22 36.84 60.94 67.41 62.80 25.51 43.98
Bi + Mean 3.90 34.56 14.27 42.71 57.89 51.15 23.26 30.56
LLM2Vec (w/o SimCSE) 24.75 49.20 39.74 65.91 69.00 67.85 25.59 48.84
LLM2Vec 39.19 53.09 41.99 78.01 71.88 75.86 31.45 56.23
Echo 12.58 49.79 36.32 68.95 70.22 67.43 26.44 45.32
Table 1: Unsupervised results on MTEB. We compare S-LLaMA-1.3B, LLaMA-2-7B, Mistral-7B, and Meta-LLaMA-3-8B with and without LLM2Vec to the unsupervised BERT models of Gao et al. (2021) as well as Echo embeddings (Springer et al., 2024).
@@
Table 1 shows the results of the best performing models, which we select based on the ablation above, on the full MTEB dataset. After the first two steps of LLM2Vec—bidirectional attention and MNTP—we observe a considerable improvement in performance for all four models (e.g., $16.4\%$ improvement for Mistral-7B).
When comparing to Echo embeddings, LLM2Vec (the first two steps only) [^3] We only directly compare the performance after the first two steps of LLM2Vec to Echo embeddings as applying SimCSE involves learning sequence representation, which makes the comparison unfair. leads to improved performance for S-LLaMA-1.3B, LLaMA-2-7B, and Meta-LLaMA-3-8B, and performs almost on par for Mistral-7B. However, compared to Echo embeddings, LLM2Vec is much more efficient as Echo embeddings repeat the input and therefore double the sequence length which makes inference considerably slower (we provide a runtime comparison in Section E.2). Adding the final step of the LLM2Vec recipe—unsupervised SimCSE—further boosts all three models by a large margin, making our LLM2Vec Mistral-7B SOTA among all unsupervised models with a score of $56.80$.
Interestingly, Meta-LLaMA-3-8B with LLM2Vec (w/o SimCSE) outperforms echo embeddings by a larger margin compared to the other models. Adding SimCSE again boosts performance, but does not outperform LLM2Vec applied to Mistral-7B.
Overall, our results highlight that LLM2Vec is successful at transforming decoder-only LLMs into strong text embedding models which outperform previous unsupervised approaches on the challenging MTEB leaderboard.